40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 49 of 53  Not logged in ELOG logo
ID Date Author Type Category Subject
  201   Wed Jul 7 20:02:53 2010 taraDailyProgressComputersBeam output is more stable

 

1) I just found out that the pbs + 1/4 waveplate in front of RefCav is not well aligned, so the length,as seen by the beam, is not 1/4 wavelength.

    Hence, the polarization is not turned 90 degree after double passing, and

    The reflected beam can go back to the laser and causes the power fluctuation we saw before.

    When the beam is blocked anywhere before RefCav, the beam output from PMC is very stable.

    I adjusted the PBS's angle and reduced the reflected power. Now the input power to PMC can go up to 50 mW without any fluctuation. 

 

2) I re-aligned the beam into RefCav, with Frank's help on gain adjusting,

    the transmitted power seems to be more stable. RefCav transmitted power RIN is posted below. I'll post the comparison between result at 40m soon. From a quick glance, RIN from 40m is at least 2 order of magnitude below

our result.

 

3) The PID control for slow actuator is up. I was adjusting it, but the medm screen was frozen.

    I reset it, and set the PID control (only P-part). The current setting for Proportional control(C3:PSL-FSS_SLOWKP) is +0.41.

 

Attachment 1: PSL_RF_RIN.png
PSL_RF_RIN.png
  200   Wed Jul 7 16:08:40 2010 MeganSummaryVCOUpdate on VCO offset

I realized I was doing the calculations incorrectly - this should make more sense.

Attachment 1: VCO_Output_Input.png
VCO_Output_Input.png
  199   Wed Jul 7 15:15:59 2010 MeganDailyProgressVCOPhase Noise Measurements

I'm currently working on measuring the phase noise of the Marconis at 80 MHz - I will be moving on to 160 MHz soon. I'm also working on learning how to make the computer do what I want it to, but I should be done with the measurements and post graphs later today. Then depending how long the measurements take today, I'll start measuring our Marconi and then move on to the VCO tomorrow and should be able to modify the VCO by Friday. One of the main things that's slowing me down is getting comfortable with processing the data on the computer.

Also, I've been having problems getting the Marconis to lock at any feedback gain below 2000. I've been using that to stay consistent and get a good lock between the two, because with lower gain there was always a sneaky little sine wave making it through the feedback loop and into the locked signal. I've accounted for this in the calibrations I've been making, with a UGF of around 1000 Hz.

  198   Wed Jul 7 10:44:39 2010 MeganSummaryVCOVCO Output vs Input

I redid one of the graphs for the output of the VCO vs the input voltage to make sure I didn't miscalculate. I got the same graph when accounting for the 19dBm attenuation at the output of the VCO. I also made a graph of the output RMS voltage ignoring the attenuation - the total output ranged from 13mV to 918mV, while the voltage added by taking the attenuation into account is 1.9929V.

I used the equation N dBm = 10*log(Vrms2/50/0.001) to solve for the voltage associated with 19 dBm.

Attachment 1: VCOOutputvsInput.png
VCOOutputvsInput.png
  197   Wed Jul 7 01:02:06 2010 taraDailyProgressLaserLaser driver is stationed on the electronic rack

I moved the laser driver to the electronic rack and it should stay there for the final setup.

A cable for interlock is made and connected to the laser driver.

A cable for slow actuator is also made and connected. Now we can use the medm FSS screen instead of the voltage calibrator

to adjust the NPRO's temperature (slow actuator.)

 

I went to 40m lab for a measurement

 

 

  196   Fri Jul 2 13:45:55 2010 FrankNotesEnvironmentchanged lab temperature

lowered the lab temperature avout 1K to get some data for the time constant from the environment to the chamber over the weekend which i can compare with my theoretical model.

current heater voltage for revcav is 14.5V @37Ohm resistance, so heater power is currently 5.68W.
current chamber temp is ~29C
current RT is 21.3C

  195   Thu Jul 1 13:49:58 2010 MeganNotesVCOVCO Output Amplitude Depending on Modlevel Input

I measured the output amplitude of the VCO with different Modlevel voltage inputs (from 0 V to 5 V) by measuring the RMS voltage on the oscilloscope. While I was varying the voltage input, I discovered the sine wave is very distorted between roughly 0.8 V and 3.5 V of modlevel input. Extra peaks started appearing that disappeared or became less prominent when below or above that range. The graphs show the output of the VCO in dBm and voltage (input impedance of 50 Ohms), accounting for the 19 dBm attenuation added to the output of the VCO.

I also started measuring the phase noise of the complete, unmodified VCO so we have a reference for the noise of the modified version when we are done with that. It is all set up right now and I should be able to start recording data on the computer soon. After we have a reference of the noise of the unmodified VCO, we will modify the VCO and see how it affects the noise level - hopefully it drops.

Attachment 1: Output_vs_Modlevel_Voltage.png
Output_vs_Modlevel_Voltage.png
  194   Wed Jun 30 23:26:27 2010 taracDailyProgressRefCavDebugging FSS servo+Locking RefCav

summary of this entry

  RefCav was not locked becase of the RFPD's input power. I haven't checked yet if it's the power supply or the cable. Now RefCav is locked.

 

Details

I was trying to lock RefCav again, but it didn't work.

So, I checked the error signal from mixer out, saw nothing.

It turned out that RF out from the PD has low voltage output, ~10mV.

I unplugged the cable and switch to another PD's cable along with its power supply, the DC increased to ~300 mV, a good sign.

Still, no error signal from Mixer Out Channel, nothing at all.

The control loop is not closed. Now 35.5 MHz LO is connected to the EOM, and to the servo card, LO input channel. The RF signal goes to the PD input.

We should be able to see the error signal from this setup.

**********The loop in the medm screen must be enabled in order to see the error signal******.

 

     EOM  -----------------BS--------------------- [ RefCav ]

        |                           |

        LO                       PD

        |                             |

        -------- X -------------

                   |

                   --->  Error signal

 

  I mix the signals from PD and LO by a Mixer [Minicurcuit ZFM 3 S+] and see something. It's a peak crossing zero at center instead of three peaks like the error signals. 

So I'm not sure what I'm seeing, but it seems that Mixer Out channel from the spare FSS servo might not working. I switched back to the first card, and now I see the error signals.

I set the RF to 6.1 V which corresponds to 97 mV of the error signal's P-P height.

Phase adj: is 4.5122 V,

phase flip: 0

Gain: [not set yet]

      All these values are saved in startup

  To summary, RefCav was not locked becase of the power cable which I haven't checked yet if it's the power supply or the cable. Now it's locked.

I'll set the gain tomorrow and see where I can connect the slow actuator for the laser driver. Right now I can't find where the servo for slow actuator would be.

 

  193   Wed Jun 30 15:51:05 2010 FrankHowToNoiseBudgetview factor for two cylinders to calculate heat transfer via radiation

view factor for two cylinders to calculate heat transfer to and from cavity via radiation

from Michael F. Modest  "Radiative Heat Transfer", Second Edition
ISBN: 0-12-503163-7

view_factor.png
 

  192   Wed Jun 30 12:39:17 2010 FrankHowToPMCPMC fixed

Quote:

Quote:

the reason why you had this flickering problem was that you had too much power on the RFPD in reflection of the PMC. You already saturated it.
I also reduced the RF power as the error signals were not signals anymore, just spikes.

my new settings are:

RF power : 3.0
Phase : 2.5
PMC Gain: 14dB

reduced laser power to 40mW. Transmitted power is 32mW .You have to exchange the output coupler mirror in front of the RFPD in order to increase the power. I think 32mW is enough, it's something like 13mW per cavity.

 

 

 Ok, I'll find another output coupler mirror and replace the current one, and make sure that it will not saturate the RFPD.

 the saturating power seems to be much to low. it saturates at .5V DC, usually you can have something like 2V or so. So we should propably fix the PD even if it is working with lower power levels.

For now it's much more important to connect all the signals to the DAQ and lock the refcav. You still have to make a lot of cables, like the ones going to the laser (fast & slow), RFPD DC, PMC and refcav transmitted light, refcav RFPD DC etc.

  191   Wed Jun 30 12:31:22 2010 taracHowToPMCPMC fixed

Quote:

the reason why you had this flickering problem was that you had too much power on the RFPD in reflection of the PMC. You already saturated it.
I also reduced the RF power as the error signals were not signals anymore, just spikes.

my new settings are:

RF power : 3.0
Phase : 2.5
PMC Gain: 14dB

reduced laser power to 40mW. Transmitted power is 32mW .You have to exchange the output coupler mirror in front of the RFPD in order to increase the power. I think 32mW is enough, it's something like 13mW per cavity.

 

 

 Ok, I'll find another output coupler mirror and replace the current one, and make sure that it will not saturate the RFPD.

  190   Wed Jun 30 11:34:37 2010 FrankHowToPMCPMC fixed

the reason why you had this flickering problem was that you had too much power on the RFPD in reflection of the PMC. You already saturated it.
I also reduced the RF power as the error signals were not signals anymore, just spikes.

my new settings are:

RF power : 3.0
Phase : 2.5
PMC Gain: 14dB

reduced laser power to 40mW. Transmitted power is 32mW .You have to exchange the output coupler mirror in front of the RFPD in order to increase the power. I think 32mW is enough, it's something like 13mW per cavity.

 

 

  189   Wed Jun 30 00:59:13 2010 FrankHowToComputersPMC servo debugging

Quote:

Quote:

 I was going to check the TF on each stage of PMC's servo.

Unfortunately, I couldn't find the floppy disc drive, so I slide the sliders (gain, RF power) around. When I add more RF power (from 1V to 7V) to 21.5 MHz EOM, the oscilaltion subsides*.

 How sad. Stop using the floppies and get one of the GPIB-Ethernet converters from Dmass. You can download the python scripts from the 40m wiki.

 we already have one but i was waiting for one the wireless bridge devices someone wanted to buy to make it wireless.

But why do you need a floppy to measure a TF?

  188   Wed Jun 30 00:37:49 2010 ranaHowToComputersPMC servo debugging

Quote:

 I was going to check the TF on each stage of PMC's servo.

Unfortunately, I couldn't find the floppy disc drive, so I slide the sliders (gain, RF power) around. When I add more RF power (from 1V to 7V) to 21.5 MHz EOM, the oscilaltion subsides*.

 How sad. Stop using the floppies and get one of the GPIB-Ethernet converters from Dmass. You can download the python scripts from the 40m wiki.

  187   Tue Jun 29 23:56:38 2010 taracNotesPMCPMC servo debugging

 I was going to check the TF on each stage of PMC's servo.

Unfortunately, I couldn't find the floppy disc drive, so I slide the sliders (gain, RF power) around. When I add more RF power (from 1V to 7V) to 21.5 MHz EOM, the oscilaltion subsides*.

(*5 mins later I came back to turn down the RF power to 1 V again, and the beam was perfectly locked for a few minutes before fluctuated again)

It's not a long term solution, but I note this for further debugging.

One thing about the gain, I see the TF of the whole PMC servo, when I increase the common gain from 0 to ~8 dB, the magnitude of the TF gets higher. If I increase more gain to ~ 10dB or something, the magnitude goes down . So there might be sth wrong about the opamp that controls the gain.

 

  186   Tue Jun 29 17:57:40 2010 taracNotesPMCcoupling eff vs Temperature

We had a problem with power coupling efficiency to the PMC.  Sometime it changes from 80% to 40% or less.

When the temperature of the NPRO is varied by a voltage calibrator @ slow actuator input of the laser controller,

at certain T(V), the coupling efficiency drops from 80% to 40% or less. This might be mode hopping.  However, the transmitted beam power still oscillates ,

and one can see the beam spot pulsating on the screen. I thought it might be mode hopping too, but after changing T, the fluctuation of the power has not gone, there is some suppression though. So it might be the servo. I'll check the TF of the servo and compare it with the schematic, D980352-B-1.

  185   Tue Jun 29 15:13:26 2010 FrankNotesRefCavrefcav physical properties

refcav physical properties from CAD model for heat capacity calculation

volume (incl. mirrors, center bore, venting hole) : 23.47900 (in^3) = 0.000384751876 m^3

mass (incl. mirrors, center bore, venting hole) : 1.86900 pounds = 0.84776414 kilograms

 

  184   Tue Jun 29 14:59:52 2010 FrankNotesNoiseBudgetbetter list of emissivity coefficients

source : http://www.omega.com/temperature/z/pdf/z088-089.pdf

emission_coefficients.pdf

  183   Tue Jun 29 14:22:31 2010 FrankNotesNoiseBudgetsummary of emissivity coefficients for refcav experiment

emissivity coefficients - ε -  (300K)

Quartz, Rough, Fused  :  0.93
Stainless Steel, sandblasted : 0.38
Stainless steel: type 18-8, sandblasted
: 0.44
Stainless Steel, polished : 0.075
Aluminum Foil, shiny side : 0.03
Electroplated Gold : 0.03

Gold: plated on stainless steel and polished : 0.028

  182   Tue Jun 29 14:20:04 2010 taracNotesLaserCalibration [Hz/V] for modulating signal (PMC and laser)

1) DC ext channel on PMC servo:          32.82 MHz/ V

2) Fast channel on the laser controller:   3.07 MHz/V ( Our Fn generator can provide +/- 20V -> 60 MHz span, this won't let us see all three peaks of the error signal from RefCav which has 35.5x2 = 71 MHz span)

3) Temperature actuator on the laser control:  220 MHz/V.  (1FSR= 714 MHz, 2x cavity length = 0.42m)

 

This measurement is done by scanning the laser frequency (1 and 3)or cavity length (2) and see the error signal.

The two sidebands are 43 MHz apart , divided that by the corresponding Voltage difference of the sidebands to get the calibration. For Temperature actuator we see use the PMC's FSR instead of the sidebands for better accuracy.

  181   Tue Jun 29 13:22:51 2010 JenneNotesNoiseBudgetlist of emissivity coefficients

Quote:
Surface Material Emissivity Coefficient
- ε -

 Source?  This is probably handy to have around, but is it from someone who actually measures these things, or is it from a random number generator with the output stuck in some table on wikipedia? 

  180   Tue Jun 29 11:53:41 2010 FrankNotesNoiseBudgetlist of emissivity coefficients

source: http://www.engineeringtoolbox.com/emissivity-coefficients-d_447.html

 

Surface Material Emissivity Coefficient
- ε -
Alloy 24ST Polished 0.9
Alumina, Flame sprayed 0.8
Aluminum Commercial sheet 0.09
Aluminum Foil 0.04
Aluminum Commercial Sheet 0.09
Aluminum Heavily Oxidized 0.2 - 0.31
Aluminum Highly Polished 0. 039 - 0.057
Aluminum Anodized 0.77
Aluminum Rough 0.07
Aluminum paint  0.27 - 0.67
Antimony, polished 0.28 - 0.31
Asbestos board  0.96
Asbestos paper 0.93 - 0.945
Asphalt 0.93
Basalt 0.72
Beryllium 0.18
Beryllium, Anodized 0.9
Bismuth, bright 0.34
Black Body Matt 1.00
Black lacquer on iron 0.875
Black Parson Optical 0.95
Black Silicone Paint 0.93
Black Epoxy Paint 0.89
Black Enamel Paint 0.80
Brass Dull Plate 0.22
Brass Rolled Plate Natural Surface 0.06
Brass Polished 0.03
Brass Oxidized 600oC 0.6
Brick, red rough 0.93
Brick, fireclay 0.75
Cadmium 0.02
Carbon, not oxidized 0.81
Carbon filament 0.77
Carbon pressed filled surface 0.98
Cast Iron, newly turned 0.44
Cast Iron, turned and heated 0.60 - 0.70
Chromium polished 0.08 - 0.36
Concrete 0.85
Concrete, rough 0.94
Concrete tiles 0.63
Cotton Cloth 0.77
Copper electroplated 0.03
Copper heated and covered with thick oxide layer 0.78
Copper Polished 0.023 - 0.052
Copper Nickel Alloy, polished 0.059
Glass smooth 0.92 - 0.94
Glass, pyrex 0.85 - 0.95
Gold pure and highly polished 0.018 - 0.035
Granite 0.45
Gypsum 0.85
Ice 0.97
Inconel X Oxidized 0.71
Iron polished 0.14 - 0.38
Iron, plate rusted red 0.61
Iron, dark gray surface 0.31
Iron, rough ingot 0.87 - 0.95
Lampblack paint 0.96
Lead pure unoxidized 0.057 - 0.075
Lead Oxidized 0.43
Lime wash 0.91
Magnesium Oxide 0.20 - 0.55
Magnesium Polished 0.07 - 0.13
Marble White 0.95
Mercury liquid 0.1
Mild Steel 0.20 - 0.32
Molybdenum polished 0.05 - 0.18
Nickel, elctroplated 0.03
Nickel, polished 0.072
Nickel, oxidized 0.59 - 0.86
Nichrome wire, bright 0.65 - 0.79
Oak, planed 0.89
Oil paints, all colors 0.92 - 0.96
Paper offset 0.55
Plaster 0.98
Platinum, polished plate 0.054 - 0.104
Porcelain, glazed 0.92
Paint 0.96
Paper 0.93
Plaster, rough 0.91
Plastics 0.91
Quarts glass 0.93
Roofing paper 0.91
Rubber, hard glossy plate 0.94
Rubber Nat Hard 0.91
Rubber Nat Soft 0.86
Sand 0.76
Sawdust 0.75
Silicon Carbide 0.83 - 0.96
Silver Polished 0.02 - 0.03
Steel Oxidized 0.79
Steel Polished 0.07
Stainless Steel, weathered 0.85
Stainless Steel, polished 0.075
Stainless Steel, type 301 0.54 - 0.63
Steel Galvanized Old 0.88
Steel Galvanized New 0.23
Tile 0.97
Tin unoxidized 0.04
Titanium polished 0.19
Tungsten polished 0.04
Tungsten aged filament 0.032 - 0.35
Water 0.95 - 0.963
Wood oak 0.91
Wrought Iron 0.94
  179   Tue Jun 29 11:45:47 2010 FrankHowToNoiseBudgetunit converter

http://www.matweb.com/tools/unitconverter.aspx

  178   Tue Jun 29 11:15:40 2010 FrankNotesFoamfoam physical properties

physical properties for the new yellow foam used for insulation (Certifoam 25):

thermal conductivity k (Btu/h ft2 °F):
@ 75° F mean :  0.200
6-Month Aged Values @ 75° F mean : 0.192
fresh (As Manufactured) : 0.11

assuming ~0.2 Btu/h ft2 °F the calculated k-value in Si units is 1.136 W/m2-K which matches the R-value given below

 

density  : 29 kg/m3

 

thermal resistance (R-value, 1in thickness)  : 0.88 K-m2/W

The conversion between SI and US units of R-value is 1 h·ft²·°F/Btu = 0.176110 K·m²/W, or 1 K·m²/W = 5.678263 h·ft²·°F/Btu.

 

 English Unit

Quantity

Multiply by

Factor

Metric Unit

inch

1

25.4

25.40

mm

foot

1

0.3048

0.30

m

in2

1

645.16

645.16

mm2

ft2

1

0.09290304

0.09

m2

ft3/min

1

0.0004719

0.00

m3/s

BTU/hr

1

0.2930711

0.29

W

BTU/hr-ft-F

1

1.729577

1.73

W/m-K

BTU/hr-ft2

1

3.1546

3.15

W/m2

BTU/hr-ft2-F

1

5.67826

5.68

W/m2-K

BTU-in/hr-ft2-F

1

0.144228

0.14

W/m-K

F-ft2-hr/BTU-in

1

6.9333466

6.93

m-K/W

F-ft2-hr/BTU

1

0.176109

0.18

m2-K/W

lbf/ft3

1

16.01846

16.02

kg/m3

lbf/ft2

1

47.88026

47.88

Pa

mph

1

0.44704

0.45

m/s

 

 


  177   Mon Jun 28 19:06:40 2010 taracDailyProgressPMCPMC acting weird

This morning, the PMC couldn't be locked to the laser. The FSS servo was disabled during that time.

When the gain/ RF power were adjusted, PMC was locked to the laser, but the coupling efficiency dropped from 80% to 40%.

I try adjusting the mirror, but it's not the alignment problem because I couldn't increase the efficiency.

So after the noise budget party with Jan and Frank, I check if the EOM works looking at the error signal from mixer out.  The laser frequency is scan at +/- 10V @ 100 Hz.

I'm not sure what is the corresponding freq span, but it must be less than 43 MHz(21.5 MHz x 2) because

three peaks of the signal could not be seen with +/- 10V span, so I use a voltage calibrator to slowly adjust the temperature and see the rest of the signals.

Nevertheless, the signals are there, so I try to lock the PMC again, and now the efficiency back to almost 80% again ( I have to re align again because of the earlier adjustment. The mixer out channel is monitored when I set the PMC gain to make sure there will be no oscillation.

I'm not sure what happen, loosen connectors, mode hopping in the laser, etc. I'll see if I can track this down , otherwise we could not have a stable locking system.

  176   Fri Jun 25 18:27:59 2010 taracLaserLasermode matching to ACav

  The new driver for AOM is working, the maximum power is ~1.3w. There's a switch for int/ext signal.

We use internal signal to drive the AOM for alignment purpose.

The mode matching for ACav is on hold. Because the1st order beam coming out of the AOM looks very elliptic.

I'm not sure if it's the result of the large beam size in the AOM or the alignment problem.

the beam might clip on the edge of AOM, the beamsize is quite large, the data sheet give 80% 1st order efficiency for 1100 um diameter spot [AOM] and I adjusted the beam size to maximize the power before. I'll try decrease the beam size and see if this reduce the elliptical shape of the beam. Once the beam size on the AOM is determined, the rest of the mode matching can be calculated.

 

  175   Thu Jun 24 09:37:16 2010 ranaLaserRefCavRefCav is locked and optimized

For the sideband power, you don't really want to make it so large. All that you really want is to make sure that the shot-noise signal is bigger than the electronics noise.

Once you guys start working on the noise budget, you will see that the input referred noise of the RFPD is equivalent to roughly 2 mA of photocurrent. So, for a reasonable

signal you should set the modulation depth (Gamma) equal to ~0.3-0.5.

Remember that we are also going to be fighting against the residual RFAM from the EOM, so its important to maximize the signal amplitude relative to the RFAM. The RFAM

signal will increase with increasing modulation depth...

  174   Wed Jun 23 23:52:48 2010 taracLaserRefCavRefCav is locked and optimized

I adjust the mode matching lenses and align the beam so that the transmitted power is ~97% of the input power. Actually I scan the beam and look at the reflected power. The reflected beam has power ~3%. The knobs on one of the periscope mirror acting weird. There's still thread (~ 4 or 5 turns, I guess) left but it seems to be very sensitive to my hand pressure when I rotate the knob, making fine adjustment rather hard .

Phase shift is adjusted by looking at the error signal. The laser is scan while the error signal from the fss servo mixer out is monitored. Feedback signals from the servo to EOM and laser must be removed.

RF power for RefCav is tuned. I assume that all sidebands' power will be reflected, and only power in the carrier will be transmitted, and to maximize the error signal's slope, we need Pcarrier/ Psideband ~ 2 [Black] .

So I measure the power of the incoming beam and adjust the RF power so that the transmitted beam's power is 1/2 of the input power. Another 1/2 of the input power will be the power of 2 sidebands that reflect back. 

 I use a photo diode to see the transmitted power and adjust the gain. The goal is to maximize the gain and have a stable transmitted power. However, the power still oscillates even when I decrease the gain, more than that and it loses lock. So I can only minimum the osicllation (You can see the beam spot pulsating on the monitor.)

All values are saved in the "startup.cmd" file.

a

  173   Wed Jun 23 03:24:30 2010 ranaMiscRefCavheater-to-can time constant

In the entry below, Frank meant to add some commentary explaining that he thinks this long time constant is anomalously long.

At the 40m, it is much shorter. The insulation there is much thinner, and the can is somewhat shorter, but I think the main effect is that the steady state temperature of this can is too low. The temperature setpoint should be set to 35-40 C.

Then the time constant can be remeasured by doing a small step (~1 deg). We can't assume linearity over 10 deg - what we care about is the small signal time constant at the quiesscent temperature. Although the heating time and the cooling time will be slightly different, they will be comparable for small steps.

If the quiesscent temperature is set to be equal to the room temperature, it is insanity. We need to count on the room acting as a conductive and convective cold load since we only have a heater for an actuator and not a cooler.

  172   Tue Jun 22 18:02:32 2010 FrankMiscRefCavheater-to-can time constant

measured step function from heater to can temperature. All four sensors show the same.

time constant ~7.5 hours, have to download and fit data for a more accurate value

StepResponse_Heater-Can.png

  171   Tue Jun 22 17:52:32 2010 FrankLab InfrastructureEnvironmenttemperature time series for the last week

temperature time series for the last couple of days measured right above the table between both cavities.

AmbMon_061710-062210.png

  170   Mon Jun 21 19:19:12 2010 taracLaserPMCProgress on PSL setup: Tuning phase shift/RF power for PMC

All SMA cables for locking RefCav are made and wired up in their places.

We decide to turn the power up to 50mW before it enters PMC. With ~80% efficiency, we get 40 mW out.

I adjusted RF power and phase shift for PMC and saved it in the init file so that next time we reboot the crate, the values will be set.

Next: I'll lock the RefCav and optimize the alignment.

  169   Sun Jun 20 02:17:06 2010 ranaElectronicsRefCavphase noise of IFR + Rubidium

see this entry

  168   Wed Jun 16 19:43:00 2010 taracLaserPMCPMC and EOM alignment

We try optimizing the PMC transmission in P wave. The maximum we can get for now is ~82%.

The 21.5 EOM is prealigned. We will do fine adjustment again when insulating cables are made.

EOM alignment means to align the polarization of the beam to match the applied electric field in the EOM. This will minimize the amplitude modulating effect.

There are 3 degrees of freedom, 2 for EOM positions, another one is the polarization of the beam.

The 35.5 EOM is also pre aligned, the signal is very small, probably because of its broadband type. I use 4395A spectrum analyzer to see the peak at 35.5 MHz, but it's really hard to tell if I minimize it or not.

I see the thermal effect on EOM crystal clearly when I adjust the EOM. After I minimize the Amplitude Modulation (AM) and left the EOM for awhile, the misalignment gradually increase.

Frank suggests that the calibration of the error signal should be done, so that we can approximate how well the temperature must be controlled to reach our noise budget. I'll think about that measurement.

As I finished my elog, the peak at 35.5 just went up. Case in point.

  167   Wed Jun 16 10:26:01 2010 FrankMiscComputerslatest medm screens from Hanford

got the latest screens from Hanford in order to avoid such problems again in the future. Someone has to search and replace the channel names for the ones we wanna use. They are currently located in the home directory on fb1. will copy them to the sun workstation asap...

  166   Wed Jun 16 00:10:36 2010 FrankComputingDAQnew channels for temp ctrl of both cavities

some new channels for the temp ctrl of the two cavities, most of them for debugging purposes only

# ACav
# Sensor1
[C3:PSL-ACAV_SENS1_VOLT]
[C3:PSL-ACAV_SENS1_KELVIN]
[C3:PSL-ACAV_SENS1_MON]
[C3:PSL-ACAV_SENS1_CAL]
# Sensor2
[C3:PSL-ACAV_SENS2_VOLT]
[C3:PSL-ACAV_SENS2_KELVIN]
[C3:PSL-ACAV_SENS2_MON]
[C3:PSL-ACAV_SENS2_CAL]
# Sensor3
[C3:PSL-ACAV_SENS3_VOLT]
[C3:PSL-ACAV_SENS3_KELVIN]
[C3:PSL-ACAV_SENS3_MON]
[C3:PSL-ACAV_SENS3_CAL]
# Sensor4
[C3:PSL-ACAV_SENS4_VOLT]
[C3:PSL-ACAV_SENS4_KELVIN]
[C3:PSL-ACAV_SENS4_MON]
[C3:PSL-ACAV_SENS4_CAL]
# Ambient Sensor1
[C3:PSL-ACAV_AMB1_VOLT]
[C3:PSL-ACAV_AMB1_KELVIN]
[C3:PSL-ACAV_AMB1_MON]
[C3:PSL-ACAV_AMB1_CAL]
# Ambient Sensor2
[C3:PSL-ACAV_AMB2_VOLT]
[C3:PSL-ACAV_AMB2_KELVIN]
[C3:PSL-ACAV_AMB2_MON]
[C3:PSL-ACAV_AMB2_CAL]
# SUM signal
[C3:PSL-ACAV_TSUM_VOLT]
[C3:PSL-ACAV_TSUM_KELVIN]
[C3:PSL-ACAV_TSUM_MON]
[C3:PSL-ACAV_TSUM_CAL]
# Stack Sensor1
[C3:PSL-ACAV_STACK_VOLT]
[C3:PSL-ACAV_STACK_KELVIN]
[C3:PSL-ACAV_STACK_MON]
[C3:PSL-ACAV_STACK_CAL]
# Servo channels
[C3:PSL-ACAV_SETPT]
[C3:PSL-ACAV_SWITCH]
[C3:PSL-ACAV_CS_MON]

# RefCav
# Sensor1
[C3:PSL-RCAV_SENS1_VOLT]
[C3:PSL-RCAV_SENS1_KELVIN]
[C3:PSL-RCAV_SENS1_MON]
[C3:PSL-RCAV_SENS1_CAL]
# Sensor2
[C3:PSL-RCAV_SENS2_VOLT]
[C3:PSL-RCAV_SENS2_KELVIN]
[C3:PSL-RCAV_SENS2_MON]
[C3:PSL-RCAV_SENS2_CAL]
# Sensor3
[C3:PSL-RCAV_SENS3_VOLT]
[C3:PSL-RCAV_SENS3_KELVIN]
[C3:PSL-RCAV_SENS3_MON]
[C3:PSL-RCAV_SENS3_CAL]
# Sensor4
[C3:PSL-RCAV_SENS4_VOLT]
[C3:PSL-RCAV_SENS4_KELVIN]
[C3:PSL-RCAV_SENS4_MON]
[C3:PSL-RCAV_SENS4_CAL]
# Ambient Sensor1
[C3:PSL-RCAV_AMB1_VOLT]
[C3:PSL-RCAV_AMB1_KELVIN]
[C3:PSL-RCAV_AMB1_MON]
[C3:PSL-RCAV_AMB1_CAL]
# Ambient Sensor2
[C3:PSL-RCAV_AMB2_VOLT]
[C3:PSL-RCAV_AMB2_KELVIN]
[C3:PSL-RCAV_AMB2_MON]
[C3:PSL-RCAV_AMB2_CAL]
# SUM signal
[C3:PSL-RCAV_TSUM_VOLT]
[C3:PSL-RCAV_TSUM_KELVIN]
[C3:PSL-RCAV_TSUM_MON]
[C3:PSL-RCAV_TSUM_CAL]
# Stack Sensor1
[C3:PSL-RCAV_STACK_VOLT]
[C3:PSL-RCAV_STACK_KELVIN]
[C3:PSL-RCAV_STACK_MON]
[C3:PSL-RCAV_STACK_CAL]
# Servo channels
[C3:PSL-RCAV_SETPT]
[C3:PSL-RCAV_SWITCH]
[C3:PSL-RCAV_CS_MON]

  165   Tue Jun 15 21:28:57 2010 FrankMiscPMCPMC servo problem fixed

found the problem why the PMC servo card wasn't working: the medm screens on the sun workstation didn't have a toggle button for the disable signal for the variable gain amplifier (AD602). this signal was set to "disabled" by default and because there was no button to enable it Tara couldn't get the servo to work. added a toggle button on the screen, PMC is now locked stable

the PMC seems to be very dirty. If we lock it in s-pol the transmitted power is only 7%, even if 60% are going into it (didn't try to optimize it). the rest is dumped in the pmc. switched back to p-pol but should think about cleaning the mirrors. we also have first contact. i think it's worth a try and we can't loose much. peter has also spare mirrors (and a spare spacer)

  164   Tue Jun 15 15:48:29 2010 FrankComputingDAQDAQ "no sync" problem fixed

fixed the "no sync" problem we had since about three weeks. It was the 100pin flat ribbon cable from the timing adapter card to the A/D card. It looks ok but has problems at one of the crimped connections. If you touch it you can toggle the connection from "OK" to "not OK" via some states in between like "OK for 5 seconds". So i replaced that cable... Now DAQ is running with 4 sensors per chamber and one temp sensor for room temp measurements. will post all the channel names later...

  163   Tue Jun 15 13:36:18 2010 taracElectronics debugging PMC servo

From the bode plot, something is not quite right. I'll debug the PMC servo. My plan is

1) Measure the TF from FP1 test to FP4 (output mon), change gain setting and see if the TF change as expected.

 *note the real TF is 20log (Vpzt/ Vin) but Vpzt ~ 50 Vmon.  Vmon is connected to Vpzt with divider circuit. To get the real TF, 20log(Vpzt/Vin), the magnitude from out TF between FP1 and out mon will be added by 20log50 = 34 dB.

2) Compare it with the calculated TF from PMC schematic 

  162   Tue Jun 15 10:14:32 2010 FrankElectronics PMC servo TF

that's not the TF of the PMC servo, it's something else. look at the gain level: -100dB. that's not more than some crosscoupling. Never trust a flat response, always think if the measured form and values make sense at all !

Take a minute and think about the form and values of the TF you expect from a servo like this. Have a look into the schematic and draw the TF shape of the individual gain stages and add them to an overall TF or use LISO to simulate it.  Then measure parts of the servo step by step in order to verify that the individual parts are working as expected.

Quote:

   The weird TF result from PMC seems to be the result of the insufficient voltage input. When I increased the swept sine voltage from 2mV to 500 mV, the result of the TF becomes as expected. See fig 1.

    Before the signal is fed back to PMC, there is a PZT notch box. It is a low pass filter. It's TF looks fine (I'll update it.)

However when the TF of the servo and the notch is measured together.

 

  161   Tue Jun 15 09:44:56 2010 taracElectronics PMC servo TF

   The weird TF result from PMC seems to be the result of the insufficient voltage input. When I increased the swept sine voltage from 2mV to 500 mV, the result of the TF becomes as expected. See fig 1.

    Before the signal is fed back to PMC, there is a PMC notch box. It is a low pass filter. It's TF looks fine (fig.2.)

However, when the TF of the servo and the notch is measured together, they look shaky.

I just read Frank's comment. I'll check the schematic of the PMC servo again.

Attachment 1: PMC2_bode.png
PMC2_bode.png
Attachment 2: notch_bode.png
notch_bode.png
Attachment 3: PMC2_bode.png
PMC2_bode.png
  160   Mon Jun 14 10:16:54 2010 taracMiscOtherweird TF from PMC servo

Since the VCO has been borrowed, I decide to work on RefCav path again. To optimize the RefCav alignment, PMC will be locked to the laser while the laser frequency is modulated by a function generator. This is where a problem comes in. The PMC servo cannot catch up with the laser. Even though I lower the modulating frequency and amplitude to 4 Hz, 1 Vpp ,which roughly corresponds to ~1MHz shift in laser frequency, the PMC servo cannot stay stable for longer than a minute. so I measure the transfer function of the PMC servo, and it does not look right, see figure1. The real magnitude which goes to the PZT has higher value, since the signal comes out of the out mon channel passes through a voltage divider.

I also measured TFs of FSS servo, both Fast and PC paths look ok. I measure TF by using swept sine measurement , the source is split by a T connector, one goes to chA of SRS785, another one goes to TF1 test (where demodulated signal goes.) The output is taken from out mon (for PMC servo), fast mon (for FSS' fast path), PC mon (for FSS's PC path), and connected to chB of SRS785.

I made sure that I switch to the right path in the medm control screen when I measure the TFs

Attachment 1: PMC_bode.png
PMC_bode.png
Attachment 2: FSS_fast_bode.png
FSS_fast_bode.png
Attachment 3: FSS_PC_bode.png
FSS_PC_bode.png
  159   Sat Jun 12 17:11:26 2010 KojiElectronicsVCOPSL VCO removed and sent to LLO

As per the request from LLO, the PSL VCO was sent to the site.
They had malfunctioning of the circuit and had no spare.

We have to figure out how we continue the work.

  158   Fri Jun 11 01:04:42 2010 taracLaserLaserscanning Acav2

From yesterday, after getting TEM 00 out of the cavity, I checked the beam if it's clipped on the edge of the hole or not. There is small light on as seen by an IR viewer. Since it seems to be very small, I'll leave it as it is for now.

There was one problem. The beam was almost on the edge of the periscope's top mirror, I decided to change the height and move the periscope , and other optics in the row, side way, since the beam was really close to the edge of the opening ( I set the beam path to the center of the hole before, so it's bit off ( 5mm, 0.2") from the cavity's natural axis), and made sure that the beam is on the center of every mirror. Then, it's 4 knob adjustment which takes me a whole day again

As of now, I got TEM00, out of the cavity. I still have to adjust the lens' position to minimize the reflected beam. Before doing so, I'll prepare a mixer, a power splitter for locking the cavity.

 

One thing about the AOM, the beam after double pass is quite elliptic. I'm not sure how to correct it, and whether it's going to be a problem or not. I'll find something to read about this.

Fri Jun 11 01:00:17 2010

  157   Thu Jun 10 00:21:15 2010 taracLaserLaserscanning ACav

I'm scanning the laser to align ACav. It's a long day of adjusting 4 knobs and 1 lens (and one periscope for a while.) 

I see higher order TEM modes at the back of the cavity, but still cannot see TEM 00 yet.

Thu Jun 10 00:20:39 2010

 

I saw TEM 00 and trying to minimize the reflected power on the PD.

I just realize that the beam path is very close to the edge of the hole (see attached.) Part of the beam might be clipped.

I'll check that with IR viewer tomorrow.

The value for Voltage Calibrator is 6.17 V.

Thu Jun 10 00:58:56 2010

 

Attachment 1: Photo_42.jpg
Photo_42.jpg
  156   Wed Jun 9 03:55:41 2010 FrankLaserLaserscanning ACav

There is no attenuator connected to the AOM. The device on the AOM (the white attenuator looking like thing) is a DC-blocker which protects the AOM. The high-power attenuators which have to be used to attenuate the high RF power in order to measure it are the black, radial heatsinked parts.

Quote:

Now I'm working on aligning the beam into ACav. I got the reflected light on the PD, and I'll scan the cavity soon.

My plan on connecting the servo:

I'll use a power splitter to split 35.5 MHz signal from "LO to SERVO" channel on the crystal frequency reference card, which is driving the 35.5 MHz EOM, to beat with the PD's signal.

If the power is too low, I might use a Marconi to beat the signal for ACav, with appropriate power level.

I also need to check which power splitters and mixers are suitable for our power output. 

The demodulated signal will be filtered by a 50 Ohms low pass filter before sent to "Servo Input" channel of the Universal PDH Servo box (D0901351.)

 

The box has two knobs that allow us to change gain and LO phase manually.

From the PDH box, the "Piezo Drive Out" will be connected to the VCO's External Modulator channel.

 

 

About AOM:

     I try to adjust the voltage of the VCO that maximize the 1st order beam from AOM. I use 5 V which is maximum on the medm control screen, but I'm not sure if it's the best or not because,

the power in the 1st order still goes up even though I reach 5V (see the attached plot.) There is an attenuator on the AOM which Frank left for me. I'll check the power that goes into the AOM and check the manual again how much power it can take. If it can take more power, I'll remove the attenuator and see if I can get more efficiency. But I'll do that after aligning ACav.

 

 

  155   Wed Jun 9 03:55:41 2010 FrankLaserLaserscanning ACav

There is no attenuator connected to the AOM. The device on the AOM (the white attenuator looking like thing) is a DC-blocker which protects the AOM. The high-power attenuators which have to be used to attenuate the high RF power in order to measure it are the black, radial heatsinked parts.

Quote:

Now I'm working on aligning the beam into ACav. I got the reflected light on the PD, and I'll scan the cavity soon.

My plan on connecting the servo:

I'll use a power splitter to split 35.5 MHz signal from "LO to SERVO" channel on the crystal frequency reference card, which is driving the 35.5 MHz EOM, to beat with the PD's signal.

If the power is too low, I might use a Marconi to beat the signal for ACav, with appropriate power level.

I also need to check which power splitters and mixers are suitable for our power output. 

The demodulated signal will be filtered by a 50 Ohms low pass filter before sent to "Servo Input" channel of the Universal PDH Servo box (D0901351.)

 

The box has two knobs that allow us to change gain and LO phase manually.

From the PDH box, the "Piezo Drive Out" will be connected to the VCO's External Modulator channel.

 

 

About AOM:

     I try to adjust the voltage of the VCO that maximize the 1st order beam from AOM. I use 5 V which is maximum on the medm control screen, but I'm not sure if it's the best or not because,

the power in the 1st order still goes up even though I reach 5V (see the attached plot.) There is an attenuator on the AOM which Frank left for me. I'll check the power that goes into the AOM and check the manual again how much power it can take. If it can take more power, I'll remove the attenuator and see if I can get more efficiency. But I'll do that after aligning ACav.

 

 

  154   Wed Jun 9 03:55:40 2010 FrankLaserLaserscanning ACav

There is no attenuator connected to the AOM. The device on the AOM (the white attenuator looking like thing) is a DC-blocker which protects the AOM. The high-power attenuators which have to be used to attenuate the high RF power in order to measure it are the black, radial heatsinked parts.

Quote:

Now I'm working on aligning the beam into ACav. I got the reflected light on the PD, and I'll scan the cavity soon.

My plan on connecting the servo:

I'll use a power splitter to split 35.5 MHz signal from "LO to SERVO" channel on the crystal frequency reference card, which is driving the 35.5 MHz EOM, to beat with the PD's signal.

If the power is too low, I might use a Marconi to beat the signal for ACav, with appropriate power level.

I also need to check which power splitters and mixers are suitable for our power output. 

The demodulated signal will be filtered by a 50 Ohms low pass filter before sent to "Servo Input" channel of the Universal PDH Servo box (D0901351.)

 

The box has two knobs that allow us to change gain and LO phase manually.

From the PDH box, the "Piezo Drive Out" will be connected to the VCO's External Modulator channel.

 

 

About AOM:

     I try to adjust the voltage of the VCO that maximize the 1st order beam from AOM. I use 5 V which is maximum on the medm control screen, but I'm not sure if it's the best or not because,

the power in the 1st order still goes up even though I reach 5V (see the attached plot.) There is an attenuator on the AOM which Frank left for me. I'll check the power that goes into the AOM and check the manual again how much power it can take. If it can take more power, I'll remove the attenuator and see if I can get more efficiency. But I'll do that after aligning ACav.

 

 

  153   Tue Jun 8 20:16:21 2010 taracLaserLaserscanning ACav

Now I'm working on aligning the beam into ACav. I got the reflected light on the PD, and I'll scan the cavity soon.

My plan on connecting the servo:

I'll use a power splitter to split 35.5 MHz signal from "LO to SERVO" channel on the crystal frequency reference card, which is driving the 35.5 MHz EOM, to beat with the PD's signal.

If the power is too low, I might use a Marconi to beat the signal for ACav, with appropriate power level.

I also need to check which power splitters and mixers are suitable for our power output. 

The demodulated signal will be filtered by a 50 Ohms low pass filter before sent to "Servo Input" channel of the Universal PDH Servo box (D0901351.)

 

The box has two knobs that allow us to change gain and LO phase manually.

From the PDH box, the "Piezo Drive Out" will be connected to the VCO's External Modulator channel.

 

 

About AOM:

     I try to adjust the voltage of the VCO that maximize the 1st order beam from AOM. I use 5 V which is maximum on the medm control screen, but I'm not sure if it's the best or not because,

the power in the 1st order still goes up even though I reach 5V (see the attached plot.) There is an attenuator on the AOM which Frank left for me. I'll check the power that goes into the AOM and check the manual again how much power it can take. If it can take more power, I'll remove the attenuator and see if I can get more efficiency. But I'll do that after aligning ACav.

 

Attachment 1: aom_eff.png
aom_eff.png
  152   Tue Jun 8 03:16:21 2010 FrankLaserLaserRefCav is locked

nice good job!

you can get the names for the channels if you click on the corresponding object (slider, button, textbox, ...) using the center mouse button. You will get a green on black box containing the full channel name...

both 35.5MHz photodetectors should be ok as we used them in the previous setup already...

ELOG V3.1.3-